# First JWST spectrum of distant activity in Long Period Comet C/2024 E1 (Wierzchos)

### v1

@@ 1 Introduction
C/2024 E1 (Wierzchos), hereafter E1, was discovered by the Catalina sky survey in March 2024 at $r_{h}=8$ au (Wierzchos et al., 2024) and will have its perihelion passage at 0.7 au in January 2026. While this will be before CI will be operating, this comet presents an excellent opportunity to study how the activity drivers of such comets evolve as they approach the Sun. Understanding the behavior of comets similar to the expected CI target is of high priority for mission success. From a mission planning point of view, it is essential to be able to predict the activity level and expected production rates close to perihelion based on distant observations. On the other hand, having a detailed understanding of the range of properties and activity patterns observed among LPCs is necessary to put CI’s measurements of an individual object into the larger context of the population. E1 was the first comet with such an orbit to be discovered at such extreme distance during the JWST-era, and since ESA selected the CI mission in 2019, giving us an excellent opportunity to learn about cometary evolution over relevant distances before the mission launches.
JWST promises to be a revolutionary facility in the study of comets (Kelley et al., 2016). As a space telescope with a wavelength range (for NIRSpec) covering the 1–5 micron infrared range it is sensitive to emission features of all major drivers of cometary activity (H2O, CO, CO2) and many of the other minor species that are expected to sublimate directly from ices (as opposed to the gas species observed in the UV/visible region, that are mostly the products of photo-dissociation in the coma; e.g., Bodewits et al. 2024). The NIRSpec integral-field-unit (IFU) mode allows not only the detection of these species, but mapping of their spatial distribution, which has revealed heterogeneous outgassing in the coma of comet 29P/Schwassmann-Wachmann 1 (Faggi et al., 2024). JWST’s sensitivity is opening a new window on distant or weak cometary activity, and it has already provided the first direct evidence for water ice sublimation driven activity in Main Belt Comets (Kelley et al., 2023; Hsieh et al., 2025) and CO2 driven activity in a Centaur (Harrington Pinto et al., 2023).
@@ 2 Observations and Data Reduction
The uncalibrated data were downloaded from the Mikulski Archive for Space Telescopes and processed using version 1.16.1 of the JWST Science Calibration Pipeline with JWST Calibration Reference Data System context file 1321. The background sky exposures were subtracted from the exposures of the comet. The four comet exposures were then combined by resampling the dithered observations to create a 53 $\times$ 51 $\times$ 941 pixel data cube. A comet spectrum was extracted using a 0.4″ radius aperture centred on the brightest point (assumed to coincide with the nucleus).
@@ 3.1 Spectrum
In Fig. 1 we show the extracted spectrum of the comet. The most prominent features are labelled: these are a broad absorption around 3 µm attributed to water ice grains, and a strong emission from CO2 at 4.3 µm. Emission from water or CO gas is notably absent. The appearance of the Fresnel peak in the centre of the water ice band at around 3.1 µm indicates that the ice is crystalline (e.g., Mastrapa et al., 2009). There are also weaker absorption features visible at 1.5, 2 and 4.5 µm that can be attributed to water ice (Mastrapa et al., 2013), which are more easily seen in Fig. 2, which shows the reflectance spectrum obtained by dividing Fig. 1 by the spectrum of the G-type solar analog star SNAP-2 (Program #1128; PI: N. Luetzgendorf; publicly available data).
@@ 4 Discussion and Conclusions
The lack of water emission at 7 au was expected, but the apparent lack of CO in this comet is surprising. Previous surveys of CO vs CO2 activity have found that distant comets (beyond 3.5 au) have CO-dominated activity (Harrington Pinto et al., 2022). JWST observations of 29P revealed both CO and CO2 at 6 au from the Sun, with a global production rate ratio of CO/CO2=70 (Faggi et al., 2024). 29P is a Centaur and in the process of transitioning to a Jupiter family comet orbit – its orbit is relatively circular near 6 au and it is observed to be constantly active, with regular large scale outbursts (Miles et al., 2016), so it has clearly been processed by activity in some way. Naively, one might expect 29P to be less likely to retain hyper-volatiles such as CO, compared with a new LPC like E1, whose activity has presumably only started very recently as it approaches the Sun. Faggi et al. (2024) also observe significant spatial structure in the emission of CO and CO2 in 29P. Their coma model indicates three distinct regions of activity with local CO-to-CO2 ratios ranging from 0 to 260, implying heterogeneity of the (near-)surface ices driving this activity. This could be explained by the evolution of 29P’s nucleus during its prolonged period of violent activity in its current orbit. In this sense the lack of structure in the outgassing in E1 makes sense – a new comet could be expected to be more uniformly active across its surface.
@@
The 3 µm absorption feature due to water ice, and in particular the Fresnel peak, is more similar in appearance to those of trans-Neptunian Objects (TNOs) and inactive Centaurs observed with JWST as part of the large sample ‘DiSCo’ programme (Pinilla-Alonso et al., 2024; Licandro et al., 2024) than that of active Centaurs like 39P (Harrington Pinto et al., 2023) or 29P (Faggi et al., 2024). The TNOs and inactive Centaurs are likely larger bodies, observed at larger distances, and show surface crystalline ice, while our IFU spatial observations show that the ice in E1 is on coma particles. The Fresnel peak is characteristic of crystalline water ice and arises when a transition between volume and surface scattering occurs, possibly suggesting that the icy coma particles are large (relative to the few micron wavelength). Detailed spectral modelling for E1 is beyond the scope of this letter, but comparison with modelling of the two water-dominated inactive Centaurs by Licandro et al. (2024) could be informative: The presence of CO2 ice in absorption in the Centaur spectra (the CO2 fundamental band at 4.27 µm) suggests that the surface is likely composed of about 50% ices (water and CO2), about 30% silicates and 10 to 20% of complex organics. CO2 and water are suggested to be in intimate mixture: if this is also the case for the (sub)-surface active layer of E1, that would be consistent with the release of water and dust grains as CO2 sublimates, and support the idea that Oort cloud comets and Centaurs formed in the same region (e.g. Nesvorný et al., 2018; Gkotsinas et al., 2024).
@@
Assuming the solar energy input is distributed between thermal re-radiation and sublimation of CO2, we find an equilibrium temperature of $T_{\mathrm{eq}}=98$ K and a CO2 production rate of $4.7\times 10^{20}$ molecules m-2 s-1 at 7 au. With our measured $Q(CO_{2})=2.546\pm 0.019\times 10^{25}$ molecules s-1, this implies a small active area, with radius $\sim$ 130 m, or 0.01% of the surface if the nucleus is as large as our radius $\sim 13.7$ km limit – this is in agreement with the relatively small $Af\rho$ measured from the ground. As the comet approaches the Sun we can expect an increase in activity levels as other areas become active and/or other species start to sublimate, although the rate of brightening varies significantly between comets (Holt et al., 2024). Deeper thermal wave propagation, seasonal effects, or removal of a mantled layer could result in a significant increase in gas activity above that predicted by extrapolation of this current production rate. We expect the Fresnel peak, and then the whole 3 µm feature, will eventually disappear from the spectrum as water ice starts to sublimate, similar to the Centaurs and comets observed closer to the Sun. Protopapa et al. (2018) showed that the physical extent of the icy grain halo decreases with decreasing heliocentric distance. NIRSpec IFU provides the first opportunity to test this at meaningfully different heliocentric distances. Whether or not CO or other highly volatile species will be seen, if they are present at depths below the surface that will eventually be heated, is more difficult to predict. Studying whether or not they are present, and also the spatial distribution within the coma if they are seen, will give important clues as to where they could be within the nucleus, which will constrain evolution models. These species could be released from amorphous water ice if there is any, or from within CO2 ice (Davidsson, 2021), so spatial correlation between ices and gas will be important to investigate. This work presents results from the first of three JWST/NIRSpec IFU observations of this comet, at distances (7, 5, 3 au) bracketing the expected start of water ice sublimation, and we look forward (in detailed follow up paper(s)) to following its evolution to test these ideas.
@@ Acknowledgements
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with JWST Director’s Discretionary Program GO-06714.
@@
For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@
All JWST is available through the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute.

### v2

@@ 1 Introduction
C/2024 E1 (Wierzchos), hereafter E1, was discovered by the Catalina sky survey in March 2024 at $r_{h}=8$ au (Wierzchos et al., 2024) and will have its perihelion passage at 0.6 au in January 2026. While this will be before CI will be operating, this comet presents an excellent opportunity to study how the activity drivers of such comets evolve as they approach the Sun. Understanding the behavior of comets similar to the expected CI target is of high priority for mission success. From a mission planning point of view, it is essential to be able to predict the activity level and expected production rates close to perihelion based on distant observations. On the other hand, having a detailed understanding of the range of properties and activity patterns observed among LPCs is necessary to put CI’s measurements of an individual object into the larger context of the population. E1 was the first comet with such an orbit to be discovered at such extreme distance during the JWST-era, and since ESA selected the CI mission in 2019, giving us an excellent opportunity to learn about cometary evolution over relevant distances before the mission launches.
JWST promises to be a revolutionary facility in the study of comets (Kelley et al., 2016). As a space telescope with a wavelength range (for NIRSpec) covering the 1–5 micron infrared range it is sensitive to emission features of all major drivers of cometary activity (H2O, CO, CO2) and many of the other minor species that are expected to sublimate directly from ices (as opposed to the gas species observed in the UV/visible region, that are mostly the products of photo-dissociation in the coma; e.g., Bodewits et al. 2024). The NIRSpec integral-field-unit (IFU) mode allows not only the detection of these species, but mapping of their spatial distribution, which has revealed heterogeneous outgassing in the coma of comet 29P/Schwassmann-Wachmann 1 (Faggi et al., 2024). JWST’s sensitivity is opening a new window on distant or weak cometary activity, and it has already provided the first direct evidence for water ice sublimation driven activity in Main Belt Comets (Kelley et al., 2023; Hsieh et al., 2025) and CO2 and methane emissions in Centaurs (Harrington Pinto et al., 2023; Pinilla-Alonso et al., 2024).
@@ 2 Observations and Data Reduction
The uncalibrated data were downloaded from the Mikulski Archive for Space Telescopes and processed using version 1.16.1 of the JWST Science Calibration Pipeline with JWST Calibration Reference Data System context file 1321. The background sky exposures were subtracted from the exposures of the comet. The four comet exposures were then combined by resampling the dithered observations to create a 53 $\times$ 51 $\times$ 941 pixel data cube. A comet spectrum was extracted using a 0.4″ ($\sim$2000 km) radius aperture centred on the brightest point (assumed to coincide with the nucleus).
@@ 3.1 Spectrum
In Fig. 1 we show the extracted spectrum of the comet. The most prominent features are labelled: these are a broad absorption around 3 µm attributed to water ice grains, and a strong emission from CO2 at 4.3 µm. Emission from water (2.69 µm) or CO gas (4.67 µm) is notably absent. The appearance of the Fresnel peak in the centre of the water ice band at around 3.1 µm indicates that the ice is crystalline (e.g., Mastrapa et al., 2009). There are also weaker absorption features visible at 1.5, 2 and 4.5 µm that can be attributed to water ice (Mastrapa et al., 2013), which are more easily seen in Fig. 2, which shows the reflectance spectrum obtained by dividing Fig. 1 by the spectrum of the G-type solar analog star SNAP-2 (Program #1128; PI: N. Luetzgendorf; publicly available data).
@@ 4 Discussion and Conclusions
Considering the gas coma, the lack of water emission at 7 au was expected, but the apparent lack of CO in this comet is surprising. Previous surveys of CO vs CO2 activity have found that distant comets (beyond 3.5 au) have CO-dominated activity (Harrington Pinto et al., 2022). JWST observations of 29P revealed both CO and CO2 at 6 au from the Sun, with a global production rate ratio of CO/CO2 = 70 (Faggi et al., 2024). 29P is a Centaur and in the process of transitioning to a Jupiter family comet orbit – its orbit is relatively circular near 6 au and it is observed to be constantly active, with regular large scale outbursts (Miles et al., 2016), so it has clearly been processed by activity in some way. Naively, one might expect 29P to be less likely to retain hyper-volatiles such as CO, compared with a new LPC like E1, whose activity has presumably only started very recently as it approaches the Sun. Faggi et al. (2024) also observe significant spatial structure in the emission of CO and CO2 in 29P. Their coma model indicates three distinct regions of activity with local CO-to-CO2 ratios ranging from 0 to 260, implying heterogeneity of the (near-)surface ices driving this activity. This could be explained by the evolution of 29P’s nucleus during its prolonged period of violent activity in its current orbit. In this sense the lack of structure in the outgassing in E1 makes sense – a new comet could be expected to be more uniformly active across its surface.
@@
Turning to the absorption features, the 3 µm water ice band, and in particular the Fresnel peak, is qualitatively similar in appearance to those of the water-dominated trans-Neptunian Objects (TNOs) and inactive Centaurs observed with JWST as part of the large sample ‘DiSCo’ programme (Pinilla-Alonso et al., 2025; Licandro et al., 2025; De Prá et al., 2025). The Fresnel peak is not seen in JWST spectra of active Centaurs like 39P (Harrington Pinto et al., 2023) or 29P (Faggi et al., 2024). The DiSCo TNOs and Centaurs are not active bodies, and are likely larger. They show surface crystalline water ice, while our IFU spatial observations show that the water ice in E1 is on coma particles. Neither the water-dominated TNOs/Centaurs nor E1 show CO ice absorption features, while other DiSCo targets do (Pinilla-Alonso et al., 2025), but, as noted above, surface CO is expected to be lost due to thermal evolution, so we cannot conclude that the comet or the TNOs/Centaurs lack CO in their interiors based on their icy coma particle or surface composition alone, and we may yet see CO emission from E1 as the comet gets closer to the Sun.
@@
The Fresnel peak is characteristic of crystalline water ice and arises when a transition between volume and surface scattering occurs, possibly suggesting that the icy coma particles are large (relative to the few micron wavelength). Detailed spectral modelling for E1 is beyond the scope of this letter, but comparison with modelling of the two water-dominated inactive Centaurs by Licandro et al. (2025) could be informative: The presence of CO2 ice in absorption in the Centaur spectra (the CO2 fundamental band at 4.27 µm) suggests that the surface is likely composed of about 50% ices (water and CO2), about 30% silicates and 10 to 20% of complex organics. CO2 and water are suggested to be in intimate mixture: if this is also the case for the (sub)-surface active layer of E1, that would be consistent with the release of water and dust grains as CO2 sublimates, and support the idea that Oort cloud comets and Centaurs formed in the same region (e.g. Nesvorný et al., 2018; Gkotsinas et al., 2024). With the caveat that we observe the coma rather than the surface, comparison between the comet and the JWST-observed Centaurs samples the effects of different nucleus size and different orbital evolution for bodies from the same part of the disc.
@@
Assuming the solar energy input is distributed between thermal re-radiation and sublimation of CO2, we find an equilibrium temperature of $T_{\mathrm{eq}}=98$ K and a CO2 production rate of $4.7\times 10^{20}$ molecules m-2 s-1 at 7 au. With our measured $Q(CO_{2})=2.546\pm 0.019\times 10^{25}$ molecules s-1, this implies a small active area, with radius $\sim$ 130 m, or 0.01% of the surface if the nucleus is as large as our radius $\sim 13.7$ km limit – this is in agreement with the relatively small $Af\rho$ measured from the ground. As the comet approaches the Sun we can expect an increase in activity levels as other areas become active and/or other species start to sublimate, although the rate of brightening varies significantly between comets (Holt et al., 2024). Deeper thermal wave propagation, seasonal effects, or removal of a mantled layer could result in a significant increase in gas activity above that predicted by extrapolation of this current production rate. We expect the Fresnel peak, and then the whole 3 µm feature, will eventually disappear from the spectrum as water ice starts to sublimate, similar to the Centaurs and comets observed closer to the Sun (Protopapa et al., 2018; Protopapa et al., 2021, 2022; Harrington Pinto et al., 2023; Faggi et al., 2024). Protopapa et al. (2018) showed that the physical extent of the icy grain halo decreases with decreasing heliocentric distance. NIRSpec IFU provides the first opportunity to test this at meaningfully different heliocentric distances. Whether or not CO or other highly volatile species will be seen, if they are present at depths below the surface that will eventually be heated, is more difficult to predict. Studying whether or not they are present, and also the spatial distribution within the coma if they are seen, will give important clues as to where they could be within the nucleus, which will constrain evolution models. These species could be released from amorphous water ice if there is any, or from within CO2 ice (Davidsson, 2021), so spatial correlation between ices and gas will be important to investigate. This work presents results from the first of three JWST/NIRSpec IFU observations of this comet, at distances (7, 5, 3 au) bracketing the expected start of water ice sublimation, and we look forward (in detailed follow up paper(s)) to following its evolution to test these ideas.
@@ Acknowledgements
We thank the anonymous referee for helpful suggestions that improved this paper. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with JWST Director’s Discretionary Program GO-06714.
@@
All JWST is available through the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, and directly via https://doi.org/10.17909/7wnt-k845.